OpenStudy (anonymous):

Solve using the multiplication principle. -18= 19y Please show all work

jimthompson5910 (jim_thompson5910):

-18= 19y -18(1/19)= 19y (1/19) -18/19 = y y = -18/19

OpenStudy (saifoo.khan):

-18= 19y Divide 19 from both sides, y = -18/19

OpenStudy (anonymous):

Answer: Stop asking so many questions and listen in class. Medal me? (No offence)

OpenStudy (anonymous):

Wow, really? How do you know the person isn't homeschooling?

OpenStudy (anonymous):

heromiles is correct

OpenStudy (anonymous):

You dont have to help me. Im am taking classes online and I have a hard time in math especially algebra. Thanx heromiles and wielder and everybody else for your help and concern.